OpenStudy (anonymous):

How would I graph f(x)=ln(x+4) and it's inverse f`(x)=e^x-4 I figured out the domain, range, and vertical aymptote but I don't understand how to graph it

OpenStudy (perl):

you want to graph the inverse?

OpenStudy (perl):

here are the two functions on the same graph https://www.desmos.com/calculator/74y7llqol0

OpenStudy (anonymous):

Why would they be crossing their axises at the -3's?

OpenStudy (perl):

lets call the inverse of f , g f(x) = ln (x + 4) g(x) = e^x - 4 if you plug in x = 0 into g(x) you get g(0) = e^0 - 4 = 1 - 4 = -3

OpenStudy (perl):

it is because of the -4, the graph y = e^x is shifted down 4 units

OpenStudy (perl):

y = e^x -> y = e^x - 4

OpenStudy (anonymous):

Is there a specific reason to use 0?

OpenStudy (perl):

you asked why does it cross the y axis at -3, the x value is zero when it crosses the y axis, thats why i plugged in x = 0